OpenStudy (anonymous):

given an adjoint matrix, compute |A| example: 1 2 2 2 0 2 -1 1 0 0 2 1 0 0 0 2

OpenStudy (anonymous):

The inverse of A is the adjoint of A divided by its determinant. Since you have a 4 x 4 matrix, you can use this fact to show that (det(A))^3 = det(adj(A)). In your case, det(Adj(A) = 8, the product of the diagonal elements. So det(A) = 2.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

\[the \sum of two numbers is 25and their difference is19.find the numbers\]